OpenStudy (anonymous):

If you know the components of a vector, what mathematical relationship can be used to find the magnitude of the vector?

OpenStudy (waynex):

Each component is a side of a right triangle. Pythagoras would give you what you want.

OpenStudy (anonymous):

elaborating on what waynex said, if you take the sqrt(a^2+b^2)=c a, and b being compoinents and c being the resultant. -cheers

OpenStudy (anonymous):

If you get a horizontal component 4x and a vertical component 3y then we can say the vector is 4x+3y. To find the magnitude all we have to do is find it's modulus, so we do the following: \[\sqrt{\left( 4^{2}+3^{2} \right)}\] Using this you will get the magnitude.

OpenStudy (anonymous):

magnitude is actually the length of a vector eg if A and B are vector with position vector a(xi+yj+zk) and b(ai+bj+ck) then the magnitude of AB is \[\sqrt{(a-x)^{2}+(b-y)^{2}+(c-z)^{2}}\] 

OpenStudy (anonymous):

\[x= a i+bj+ck\] then, \[|x|=\sqrt{a^{2}+b^{2}+c^{2}}\]